# Over-Tokenized Transformer: Vocabulary is Generally Worth Scaling

### v1

@@ Abstract
Machine Learning, Large Language Modeling, Tokenization, Pre-training, Scaling-laws
@@ 2.1 Tokenization Design
Tokenization is a critical component in the development of large language models (LLMs). Established methods such as Byte-Pair Encoding (BPE) (Sennrich, 2015) and Unigram Language Models (Kudo, 2018) have been widely adopted to create subword vocabularies that balance sequence length and vocabulary size. Recently, alternative paradigms have been proposed to address the computational inefficiencies of byte-level models (Xue et al., 2022; Yu et al., 2023). For instance, MegaByte (Yu et al., 2023) combines byte-level modeling with patching by first predicting patches and then predicting bytes within each patch, improving processing efficiency. However, its performance at scale remains inferior to that of tokenizer-based approaches. Building on this concept, In this work, our findings suggest that applying n-gram patching on top of BPE expands the vocabulary size, improving model capability, while BLT (Pagnoni et al., 2024) also has the same finding on byte-level models.
@@ 2.2 Scaling Vocabulary
Recent empirical studies have systematically investigated the relationship between vocabulary size and model performance.  (Tao et al., 2024) demonstrates that expanded vocabularies enhance both training efficiency and model performance, particularly in larger architectures. Building on these findings, we argue that vocabulary size research should separately consider embedding (input) and unembedding (output). While embedding incurs only lookup costs, unembedding introduces computational costs that scale with vocabulary size. More importantly, we find that input and output vocabularies exhibit distinct scaling behaviors, underscoring the need for independent scaling strategies to optimize model design.
2.3 Multi-Token Prediction and n-Gram Modeling
Multi-Token Prediction (MTP) (Gloeckle et al., 2024) has advanced the field of next-token prediction through the introduction of auxiliary objectives for simultaneous multi-token prediction. This methodology shares fundamental principles with our $n$-gram modeling framework, where multi-token prediction can be theoretically formulated as an approximation of $n$-gram unembedding utilization (over-decoded models). In our work, we further compare the effectiveness of MTP and over-encoded models, demonstrating that the performance gains from MTP and over-encoded models are complementary and can be combined to achieve even greater improvements.
@@ 3.2 Over-Tokenized Transformer
$\texttt{OE}(x)=\mathbb{E}^{V\times d}(x^{(-1)})+\sum_{i=2}^{n}\mathbb{E}^{m\times\frac{d}{n}|k}(x^{(-n)}),$ (5)
@@ B.4 OLMoE Experiments
We present detailed metrics comparing OT-12.8M and OE-12.8M on OLMoE-1.3B in Figure LABEL:olmoe_1b3_ot_all_metrics. OT is not always better than OE, e.g., Winogrande and Social IQA.

### v2

@@ Abstract
Machine Learning, ICML
@@ 2.1 Tokenization Design
Tokenization is a critical component in the development of large language models (LLMs). Established methods such as Byte-Pair Encoding (BPE) (Sennrich, 2015) and Unigram Language Models (Kudo, 2018) have been widely adopted to create subword vocabularies that balance sequence length and vocabulary size. Recently, alternative paradigms have been proposed to address the computational inefficiencies of byte-level models (Xue et al., 2022; Clark et al., 2022; Tay et al., 2022; Yu et al., 2023). For instance, MegaByte (Yu et al., 2023) combines byte-level modeling with patching by first predicting patches and then predicting bytes within each patch, improving processing efficiency. However, its performance at scale remains inferior to that of tokenizer-based approaches. Building on this concept, In this work, our findings suggest that applying $n$-gram embeddings on top of BPE expands the vocabulary size, improving model capability, while BLT (Pagnoni et al., 2024) also has the same finding on byte-level models.
@@ 2.2 Scaling Vocabulary
Huang et al. (2021) firstly explores recurrent neural networks with scalable $n$-gram embedding table, where modular hashing is used to control vocabulary size. Roy et al. (2022) further augments the Transformer architecture with $n$-grams that are constructed from a discrete latent representation of the text sequence. Recent empirical studies have systematically investigated the relationship between vocabulary size and model performance. Tao et al. (2024) demonstrates that expanded vocabularies enhance both training efficiency and model performance, particularly in larger architectures. Building on these findings, we argue that vocabulary size research should separately consider embedding (input) and unembedding (output). While embedding incurs only lookup costs, unembedding introduces computational costs that scale with vocabulary size. More importantly, we find that input and output vocabularies exhibit distinct scaling behaviors, underscoring the need for independent scaling strategies to optimize model design. Moreover, Liu et al. (2024) revisits $n$-gram language modeling at massive scale, proposing an $\infty$-gram model trained on 5 trillion tokens, which achieves competitive performance and can effectively complement neural LMs to reduce perplexity.
2.3 Multi-Token Prediction
Multi-Token Prediction (MTP) (Stern et al., 2018; Qi et al., 2020; Gloeckle et al., 2024) has advanced the field of next-token prediction through the introduction of auxiliary objectives for simultaneous multi-token prediction. This methodology shares fundamental principles with our $n$-gram modeling framework, where multi-token prediction can be theoretically formulated as an approximation of $n$-gram unembedding utilization (over-decoded models). In our work, we further compare the effectiveness of MTP and over-encoded models, demonstrating that the performance gains from MTP and over-encoded models are complementary and can be combined to achieve even greater improvements.
@@ 3.2 Over-Tokenized Transformer
$\texttt{OE}(x)=\mathbb{E}^{V\times d}(x^{(-1)})+\sum_{i=2}^{n}\mathbb{E}^{m\times\frac{d}{n}|k}(x^{(-i)}),$ (5)
@@ 4.3 Over-Tokenized Transformer
4.4 Speed Analysis
We analyze training and inference speed for over-encoding in this section. To illustrate training efficiency, we show training throughputs in OLMoE experiments in Table 6, where we run OE and baseline under the same hardware configurations. OE in OLMoE-7B yields more overhead as we did not carefully optimize engineering configurations.
OLMoE-1.3B OLMoE-7B
Hardware 32$\times$×A100 64$\times$×A100
Baseline 1.211 0.494
+OE 12.8M 1.155 -4.63% 0.453 -8.3%
Table 6: Training throughputs for OE and baseline. We report average tokens per second in millions.
Batch Size Phase Dense-1B MoE-1B/7B Dense-7B
Baseline OE-12.8M Baseline OE-12.8M Baseline OE-12.8M
1 Prefill 20728.7 19446.4 6303.8 6189.0 6571.0 6499.9
Decode 136.2 126.6 28.2 27.9 65.1 63.3
8 Prefill 36907.5 35902.6 22297.3 22292.1 - -
Decode 797.2 760.9 184.9 181.0 232.1 228.6
64 Decode 1422.1 1407.4 860.3 826.7 - -
Table 7: Inference speed for OE and baseline. The sequence length is fixed to 2048, and we report tokens per second for prefilling and decoding separately under various batchsizes. Settings that cause OOM are leave blank.
OLMoE-1.3B OLMoE-7B
baseline 0.5409 2.3578
+OE 12.8M 0.5430 +0.38% 2.3662 +0.35%
Table 8: GFLOPs per token in the forward pass.
Theoretically, the additional flops introduced by OE is less than 0.5% (as shown in the table below), as demonstrated in Table 8. The overhead measured in the experiments should mainly be introduced by the all-to-all communication (which is proportional to the size of data parallel). And these communication overheads can be further optimized through engineering techniques in the future.
For inference efficiency, we test the prefill and decoding throughput on a single A100 GPU using the transformers library. For the OE models, the additional embedding parameters are offloaded to the CPU, incurring no GPU memory overhead. The numeric results are shown in Table 7. The impact of OE on inference throughput is negligible, especially for larger models or larger batch sizes. In contrast, the sparse parameters introduced by MoE face severe memory access bottlenecks during inference. A very large batch size is required for the MoE model to achieve the same throughput as a dense model with the same activated parameters. Considering that the model inference might be carried out on more cost-effective but less computationally powerful inference GPUs, the relative overhead of OE could be further reduced.
@@ 5 Conclusion
Acknowledgements
This research was conducted at ByteDance Inc. We are grateful for the suggestions and discussions from Jundong Zhou, Zihao Huang, Yu Bao, Yike Yuan, Sijun Zhang and other colleagues at ByteDance Seed.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ B.4 OLMoE Experiments
We present detailed metrics comparing OT-12.8M and OE-12.8M on OLMoE-1.3B in Figure 13. OT is not always better than OE, e.g., Winogrande and Social IQA.
B.5 In-house Experiments
We also perform in-house experiments to verify the effectiveness of over-encoding. The in-house baseline adopts MoE architecture, having 400M activated parameters and 4B sparse parameters in total. Then we implement OE, scaling up the vocabulary size to $m=36\mathrm{M}$. We train 600B tokens with in-house pre-training datasets, and evaluate the model through a series of open benchmarks under the few-shot settings. Results are shown in Table 9. OE shows significant gains across different domains, including reasoning, knowledge and math ability. Especially, we notice that OE mostly improves knowledge ability. These results demonstrate that OE also improves in-context learning.
Table 9: Comparison of OE and baseline on in-house models. The models are evaluated on various benchmarks on the few-shot settings. OE improves both reasoning, knowledge and math related benchmarks.
Reasoning Related Benchmarks
ARC Challenge Drop BBH WinoGrande Hellaswag
Baseline 65.7 34.4 37.1 63.2 66.2
+OE 36M 67.9 36.3 39.5 65.5 67.2
Knowledge Related Benchmarks
MMLU C-Eval TriviaQA MMLU-Pro AGIEval
Baseline 54.8 61.3 39.7 21.1 39.1
+OE 36M 57.9 68.3 49.0 24.1 43.2
Math Related Benchmarks
Ape210K GSM8K MATH
Baseline 63.7 40.6 22.7
+OE 36M 63.8 46.2 25.3